• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name P. King

    Log entry time 00:04:25 on September13,2005

    Entry number 154550

    keyword=Swing shift summary

    Executive summary:
    ==================

    Attempted to run at a beam current of 40 uA, but the right septum temperature increased enough that we had to turn down the beam current to about 38 uA.

    Completed slug 77 (IHWP in) with 212k pairs in runs 4998 through 5005.

    Began slug 78 (IHWP out) with run 5006.

    The Hall A laser lost lock in the middle of the shift, but the operations crew was able to restore it. If it fails overnight, the injector on-call will not be contacted before morning.


    Details of the shift:
    =====================

    16:00 Run 4997 (Sweeper was off).

    16:25 Beam off. Move to empty for compton tune.

    16:33 Target is at empty position; Compton tuning begins.

    17:22 Compton tuning is over; move the He target into the beam.

    17:30 Helium target is in position, request 20 uA.

    17:31 Request 40 uA.

    17:36 Begin Parity run 4998.

    17:50 Bryan comes down to tell us to turn on the beam modulation.
    He requests beam modulation to be enabled.

    18:10 Beam is off for 5 minutes.

    18:13 Begin Compton run 12148.

    18:32 Begin Parity run 4999.

    18:52 Request a reduction in our beam current of about 2 uA;
    the right septum temperature had climbed to 7.636.

    18:58 MCC takes the beam to check the viewers for Hall C.

    19:08 MCC is sending beam back.

    19:10 Begin Parity run 5000; there is no beam in this run.

    19:13 Begin Parity run 5001.

    19:19 There have been and will be a lot of beam trips; these
    are due to the checkout of the ion chambers in Hall C.

    20:07 Begin Parity run 5002.

    20:19 Problems ending Compton run 12148. Unsuccessful start of
    Compton runs 12149 and 12150. Do CODA reboot for the
    Compton. Begin Compton run 12151.

    20:48 MCC has finished with the ion chamber checkout for Hall C.
    There are now ion chamber trips from the eP chamber; MCC
    requests permission to turn off the beam modulation to
    diagnose the problem: the Hall A laser has noise and is
    losing lock. Injector on-call is on the way; beam delivery
    will continue until they arrive.

    21:05 Begin Parity run 5003.

    21:30 CORRECTION: injector on-call will not be called in to fix
    the laser tonight. Beam delivery will continue as long as
    the laser works; if it fails, then Hall A will just be down
    until morning. So far, it seems to be working.

    22:02 Begin Parity run 5004.
    Begin Compton run 12152.

    22:58 Begin Parity run 5005.
    Begin Compton run 12153. There were communication errors
    with ROC0 at the end of Compton run 12152.

    23:11 Request the IHWP be taken OUT.
    Run flipper; increment slug number to 78.
    Begin Parity run 5006.
    Begin Compton run 12154.